OpenStudy (anonymous):

Which choice below is a boxplot for the following distribution? 56, 52, 93, 50, 72, 62, 45, 72, 70, 58, 66, 91, 48, 74, 77, 71, 46

OpenStudy (anonymous):

SOO confusedddddd

OpenStudy (anonymous):

those are the boxplots to chose from!!

OpenStudy (anonymous):

how do you do it??

OpenStudy (texaschic101):

45,46,48,50,52,56,58,62,66,70,71,72,72,74,77,91,93 median is 66 Q1 = (50 + 52)/2 = 102/2 = 51 Q3 = (72 + 74)/2 = 146/2 = 73 minimum value is 45 maximum value is 93 I want to say D

OpenStudy (anonymous):

you are correct! thank you!

OpenStudy (texaschic101):

your welcome :) would you like me to explain it to you ?

OpenStudy (anonymous):

what about this one pleaseeee. @texaschic101

OpenStudy (texaschic101):

first we will put the numbers in order... 2,3,3,4,4,6,7,7,8,9,10,10,11,15,17,18,19 find the median (the middle number) this number is also Q2(quartike 2)...the median is 8 do you understand so far ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is it 0?

OpenStudy (anonymous):

(the first quartile)

OpenStudy (texaschic101):

now, to find Q1(quartile 1)....use the numbers before the median... 2,3,3,4,4,6,7,7....take the middle number of these numbers. If there are two middle numbers, take the average of those two numbers....(4 + 4)/2 = 8/2 = 4 Q1 = 4 Haven't lost you yet have I ?

OpenStudy (anonymous):

ok, no i am following

OpenStudy (texaschic101):

To find Q3, take the numbers after the median... 9,10,10,11,15,17,18,19 Take the median of this set, which there is 2, so we find the average of the 2 middle numbers..(11 + 15)/2 = 26/2 = 13 Q3 = 13 now the minimum value is the lowest number....9 maximum value is the highest number....19

OpenStudy (anonymous):

yes, i dont get how you know which box it is, if there are no numbers

OpenStudy (texaschic101):

sorry....I messed up....minimum value is 2 and maximum value is 19

OpenStudy (anonymous):

ok

OpenStudy (texaschic101):

it is tricky if there is no numbers.... minimum value 2 ---->Q1 = 4...so we know that Q1 is not far from the start. median is 8 and Q3 = 13. From this, I can tell that the median will be closer to the end. And since the maximum value is 19....its not far from the end of the box. You want to take a guess at which is the right one ?

OpenStudy (texaschic101):

oops...wait...the median is closer to the middle of the box

OpenStudy (anonymous):

E?

OpenStudy (anonymous):

since the median is closer to the middle

OpenStudy (texaschic101):

I am wanting to say B.......minimum value is 2 and Q1 = 4, so the box is not gonna be far from the start of the line. The median would be a little less then halfway.

OpenStudy (anonymous):

correct! thank you! i have one more pleasee @texaschic101 ! your a hugee help!

OpenStudy (anonymous):

im stuck.. cuz its an even amount of numbers

OpenStudy (texaschic101):

start by putting numbers in order... 5,25,33,34,34,37,37,40,42,45,45,46,46,49,73 (Q2)median is 40 (Q1) = 5,25,33,34,34,37,37 = 34 (Q3) = 42,45,45,46,46,49,73 = 46 minimum value is 5 maximum value is 73 Not 100% sure on this one.....it is either C or D....the median is supposed to be in the VERY MIDDLE. But there is a quite a bit of room from Q3 to the maximum value....I keep wanting to say C, but I am not sure

OpenStudy (anonymous):

wouldnt the box be bigger?

OpenStudy (anonymous):

cuz its bigger numbers

OpenStudy (texaschic101):

Your probably right.....Q1 = 34, Q2 = 40, Q3 = 46 the box is 12 spaces......I would go with D.

OpenStudy (texaschic101):

But Q3 is a pretty good ways from the maximum value......I am not totally sure on this one

OpenStudy (anonymous):

aww.. you were right the first time! its c :[

OpenStudy (texaschic101):

oh...I am sorry

OpenStudy (anonymous):

no worries! thanks for your help! ill prob need some more help soon if you dont mind :)

OpenStudy (texaschic101):

If I am on, I will be happy to help you :)

OpenStudy (anonymous):

@texaschic101